OpenStudy (anonymous):

1 2 -1 | -1 2 5 -1 | -4 -3 -9 -5 | -5 a)R_2=-2(r1)+(r2) b)R_3=3(r1)+(r3) c)R_3=3(r2)+(r3) I got... 1 2 -1 | -1 0 1 1 | -2 0 0 -5 | ? I cannot figure out the last one ;/ 0 0 -5 0 0

OpenStudy (anonymous):

ignore those nmbers at the end please

OpenStudy (anonymous):

i think there is a mistake at row 3 check it out ?

OpenStudy (anonymous):

its all right but i cant get the last one.

OpenStudy (anonymous):

w8 i m mistaken hold on

hartnn (hartnn):

or did u get correct till here (after 2 steps you mentioned) 1 2 -1 | 1 0 1 1 |-2 0 -3 -8|-8

OpenStudy (anonymous):

yeah

hartnn (hartnn):

3 R2+R3 will give you 3(-2)+-8 = -6-8 = -14 on the right.

OpenStudy (anonymous):